OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

Yes.

OpenStudy (texaschic101):

2x + 3y = 1200 3y = -2x + 1200 y = -2/3x + 400 slope is -2/3 and y intercept is 400 as for putting it in slope intercept form, I am not sure, so this is a guess... I will be using (2,3) x1 = 2 and y1 = 3 slope(m) = -2/3 y - y1 = m(x - x1) y - 3 = -2/3(x - 2) you might need to get a second opinion on this just to make sure